three different language books, two different social science books and five mathematics books are to be arranged on a shelf. How many different arrangements are possible if the book in each group must all stand together ?
Answer by stanbon(75887)

three different language books, two different social science books and five mathematics books are to be arranged on a shelf. How many different arrangements are possible if the book in each group must all stand together ?
----
1st: arrange the groups: 3! = 6 ways
---
2nd: arrange the books in each group: 3!*2!*5! = 6*2*120 = 1440 ways
----
Total arrangements: 6*1440 = 8640
===============
Cheers,
Stan H.